Let us consider a market where 3 firms I = {1, 2, 3} compete `a la...

  1. Let us consider a market where 3 firms I = {1, 2, 3} compete `a la Cournot (quantity-setting competition). The inverse demand function is given by p(Q) = 300 − 5Q, where Q = q1 + q2 + q3. The cost function is homogeneous and it is C1(q) = C2(q) = C3(q) = 30q.

    • Write explicitly the profit functions of each i ∈ I.

    • Derive best reply functions and the Nash equilibrium of the game.
Answer #1

P= 300 - 50 P = 300 - 5121+22+ 23) From = = PEI- 3091 [ 300 - 5(21+92 +93)] 2, - 309, 30001 - 5121+22+23) &, - 302, = 2702, -54 - 2,- qz = 222 q2 = 54-91-23 firmizs best response ph q2 = 54-91-93 in Similarly firm 3. 53= P23 - 3093 = [ 300 - 51917 9&, = 54 - 22-93 2 29,= 54-92-93 29, +q2t 93 = 54 92 = 54 - 21-93 2 292 = 54-2,-23 21+2q2 +93 = 54 23= 54-21-92 283= 54-{1-92TAL = 54 54 2 54 1 2 = 54(4-1) - 1(108 - 54) +1 (54-108) = 162 - 54-54 = 54 A2 = 2 54 1 54 ) = 1 54 2 2.0108 - 54) - 54 (2-1)